OpenStudy (anonymous):

A stone is dropped in a lake, creatinga circular ripple that travels outward at a speed of 60cm/sec. Express the area of this circle as a function of time t(in seconds)

OpenStudy (anonymous):

if it is moving with constant velocity then x(t)=60*t here x(t) is the radius of circle so A=pie*r^2 so =3600t^2*pie

OpenStudy (anonymous):

Thank you! I have another if you're interested...

OpenStudy (anonymous):

yes you may give

OpenStudy (anonymous):

An airplane is flying at a speed of 350 mi/h. at an altitude of one mile. the plane passes directly above a radar station at time t=0...A) Express the distance s (in miles) between the plane and the radar station as a function of the horizontal distance d (in miles0 that the plane has flown....B) Express d as a function of the time t (in hours) that the plane has flown...